# PiVe: Prompting with Iterative Verification Improving Graph-based Generative Capability of LLMs

### v2

@@ Abstract
Large language models (LLMs) have shown great abilities of solving various natural language tasks in different domains. Due to the training objective of LLMs and their pre-training data, LLMs are not very well equipped for tasks involving structured data generation. We propose a framework, Prompting with Iterative Verification (PiVe), to improve graph-based generative capability of LLMs. We show how a small language model could be trained to act as a verifier module for the output of an LLM (i.e., ChatGPT, GPT-4), and to iteratively improve its performance via fine-grained corrective instructions. We also show how the verifier module could apply iterative corrections offline for a more cost-effective solution to the text-to-graph generation task. Experiments on three graph-based datasets show consistent improvement gained via PiVe. Additionally, we create GenWiki-HIQ and highlight that the verifier module can be used as a data augmentation tool to help improve the quality of automatically generated parallel text-graph datasets. [^1] Our code is at https://github.com/Jiuzhouh/PiVe.
@@ 1 Introduction
A semantic graphs, as a form of graph-structured data, stores information in a machine-accessible way (van Harmelen et al., 2008). Generating a semantic graph from text is known as text-to-graph (T2G) generation and is previously attempted mostly by fine-tuning small language models (Xu et al., 2022; Guo et al., 2020). However, generating graph-structured data remains a challenge for LLMs even in the presence of reasonable number of few-shot examples. In fact, regardless of the number of few-shot examples or prompting style the outputs from LLMs (e.g., GPT-3.5) still contain errors and require correction (§4.4).
@@ 2 Basic Definitions
To prompt the LLM, we use few-shot by showing an example of T2G in the prompt to specify the expected format of the semantic graph (i.e., set of triples). We report experiments under various number of shots (§4.6). The basic form of instruction we use in the prompt is "Transform the text into a semantic graph." followed by a text and a semantic graph pair as a demonstration. We also show results under different prompting strategies (§4.7). Different demonstrations are used for different datasets to adapt to the style of different datasets. We show the used demonstrations in Appendix I.
@@ 3.2 Iterative Prompting
As with the prompt used in the first iteration, we still provide an example in the new prompt for the subsequent iterations. The new prompt is "Transform the text into a semantic graph and also add the given triples to the generated semantic graph." In addition to the text, we also include some triples predicted by the verifier module which LLMs are likely to miss. This explicitly instructs the LLM to generate semantic graph and include the given triples. The given triple set contains the predicted missing triples from each iteration, which prevents the LLM from making the same mistakes as in previous iterations. See Appendix I for the used demonstrations.
@@ 4.2 The LLM and Verifier Modules
ChatGPT (gpt-3.5-turbo) is used as our default LLM to perform the T2G task. [^3] We also compare the graph-based generative capability between ChatGPT and GPT-3 in Appendix D and report the fine-tuned results on small language model in Appendix H. We also experiment with GPT-4 in Subsection 4.7. For verifier module, we use T5-Large (Raffel et al., 2020), and Flan-T5-XXL (Chung et al., 2022) as the backbone models for dataset-specific verifier module, and unified verifier module, respectively. T5 models follow the encoder-decoder architecture and treat all NLP tasks as unified text-to-text transduction tasks. Flan-T5 is instruction-fine-tuned version of T5 which was trained on 1,836 NLP tasks initialized from fine-tuned T5 checkpoint. For T5-large, we fine-tune all parameters for separate verifier modules per each dataset. While for Flan-T5-XXL, we use LoRA (Hu et al., 2022) as a parameter-efficient fine-tuning method, to train a unified verifier module which can follow the instruction. When using the unified verifier, we specify the dataset name in the instructions as datasets have different naming convention for relations.
The verifier are implemented using Pytorch (Paszke et al., 2019) and Transformers (Wolf et al., 2020). For the training, we use Adam optimizer (Kingma and Ba, 2015). Details about hyperparameter setting is provided in Appendix A. For the implementation of parameter efficient training method used in Flan-T5-XXL, we use PEFT (Mangrulkar et al., 2022) and 8-bit quantization technique (Dettmers et al., 2022). All training was done using a single A40 GPU with 48GB of RAM.
@@ 4.3 Evaluation Metrics
T-F1↑ G-F1↑ G-BS↑ GED↓ T-F1↑ G-F1↑ G-BS↑ GED↓
Single Verifier Base 13.50 4.89 83.92 13.20 13.50 4.89 83.92 13.20
Iteration 1 17.92 5.78 85.91 12.37 17.76 5.83 86.42 12.37
Iteration 2 19.46 6.44 86.57 12.08 18.51 6.11 86.91 12.19
Iteration 3 20.17 6.61 86.83 11.95 18.55 6.17 86.94 12.18
Unified Verifier Base 13.50 4.89 83.92 13.20 13.50 4.89 83.92 13.20
Iteration 1 19.64 6.00 86.39 12.08 16.99 5.67 87.08 12.95
Iteration 2 22.11 6.44 87.31 11.68 17.76 5.67 87.48 12.96
Iteration 3 23.11 7.50 87.70 11.35 17.85 5.67 87.52 12.96
Table 3: Comparison between Iterative Prompting and Iterative Offline Correction on KELM-sub dataset across all metrics using Single Verifier and Unified Verifier.
@@ 4.4 Results
On WebNLG and GenWiki datasets, single verifier module performs slightly better than unified verifier module. While on KELM-sub dataset, unified module performs far better. We speculate this is due to the size of training data for KELM-sub verifier module being larger than that for WebNLG and GenWiki (as shown in Table 1). Since unified verifier module combines the training data of different datasets, more training data leads to better performance for instruction-tuning. We demonstrate two qualitative examples in Appendix G and also conducted human evaluation which we include in Appendix F due to page limit.
@@
Instead of iteratively prompting the LLM, another way to utilise the results from verifier module is to append the predicted missing triples to the previously generated graph. The results of the comparison between iterative prompting and iterative offline correction using single verifier module and unified verifier module on KELM dataset is shown in Table 3. Iterative offline correction performs worse than iteratively prompting. This might be because iteratively prompting has the chance of doing self-correction. In each iteration, when we prompt the LLM, the generated graphs can probably correct the mistakes that were made in previous iteration. For example, in Figure 6 of the Appendix, in Base the predicted relation regarding birth date is “birth year”, while the reference is “date of birth”. As the PiVe iteration continues, in Iteration 2, the relation “birth year” is regenerated as “date of birth” even though we didn’t mention this in the prompt. Due to the page limit, we report the comparison results on WebNLG and GenWiki datasets in Appendix B. Similarly, iterative prompting can achieve better results than iterative offline correction over all using different verifier modules.
@@ 4.6 Impact of More Shots
While in our main experiments, for cost reason, we used only one-shot demonstrations for the LLM prompting (i.e., GPT-3.5), we show that PiVe is effective in improving the results regardless of the underlying number of shots. Here we report the results of k-shot (k=6, 8, 10) with the iterative offline correction (i.e., only using the LLM once to get the initial graph, while the correction steps are all applied step-by-step and offline). Figure 2 demonstrates the results on KELM-sub using unified verifier with iterative offline correction (for detailed numbers see Table 10 in Appendix). The results show, as expected, that PiVe still provides consistent improvement even with the increase in the number of shots. Additionally, as the shots grow the improvement from PiVe also increases.
@@ 4.7 Baselines on LLMs
Figure 3: Results of using 3 diverse prompts with 6-shot on KELM-sub with Iterative Offline Correction on ChatGPT and GPT-4. The colors represent Base, and corrective iterations 1, 2, 3. Prompt 1: Transform the text into a semantic graph. Prompt 2: Transform the text into a semantic graph consisting of a set of triples. Generate as many triples as possible. Prompt 3: Transform the text into a semantic graph consisting of a set of triples. First produce all relations possible, then produce the graph.
@@ 4.8 Computational Cost and Trade-off
The training and inference of both single verifiers and unified verifier are on a single A40 GPU. Each single verifier takes around 6 hours and the unified verifier takes around 40 hours to train. The computation cost for training of verifiers is a feasible one-off cost. Once the training is finished, the inference of the verification of each instance takes 0.15s for single verifier, and 3.5s for unified verifier. Different verifiers present performance-speed trade-offs and are significantly effective in augmenting the LLMs.
@@ 5 GenWiki-HIQ
Table 4: Results of iterative augmentation on filtered 110K text-graph pairs from GenWikiFINE across four metrics after each iteration. We take the text-graph pairs from Iteration 4 as the created GenWiki-HIQ dataset.
@@
We used the dataset-specific verifier module to do the data augmentation. We conducted four iterations and the evaluation results are shown in Table 4. The results in Base represent the scores over non-parallel graph-text pairs from GenWikiFINE, which have low overlap between graph and text. By using verifier module iteratively, we add more missing triples to the original graph, thus leading the higher quality scores. As the iteration progresses, fewer missing triples are added and we take the augmented graph-text pairs from the last iteration as the final created GenWiki-HIQ dataset. We also conducted G2T experiments in Appendix E to further demonstrate the quality of GenWiki-HIQ. The G2T model trained on GenWiki-HIQ performs far better than the G2T model trained on GenWikiFINE-f on the human annotated GenWiki test set. This indicates that GenWiki-HIQ contains parallel text-graph pairs with high overlap.
@@ 6.3 Verifiers
Leveraging small models could further improve the performance of LLMs. Cobbe et al. (2021) proposed to solve math word problem by utilising verifier. The verifier is used to judge the correctness of model-generated solutions. During test time, based on multiple candidate solutions generated, verifier calculates the correctness probability and the final answer will be selected by the verifier from the ranked list. Welleck et al. (2023) proposed self-correction, an approach that trains a small models to iteratively apply self-correction. The idea of self-correction looks similar to our PiVe. While Welleck et al. (2023) focuses on the design of a self-correcting language model, PiVe presents a very simple verifier module design and a simple data perturbation strategy to train such model. The ideas presented in our work are developed concurrently and independently.
@@ Appendix B Additional Experiment Result
6-shot Base 35.25 10.78 85.43 10.19
Iteration 1 36.96 12.56 88.20 10.14
Iteration 2 37.25 12.61 88.47 10.13
Iteration 3 37.37 12.68 88.54 10.13
8-shot Base 36.40 14.89 86.63 9.59
Iteration 1 39.51 17.78 89.24 8.86
Iteration 2 39.80 18.11 89.38 8.81
Iteration 3 39.82 18.17 89.38 8.81
10-shot Base 37.14 14.72 86.57 9.46
Iteration 1 40.40 18.61 89.46 8.64
Iteration 2 40.69 19.28 89.66 8.56
Iteration 3 40.70 19.33 89.68 8.55
Table 10: Results of doing k-shot (k=6, 8, 10) learning on KELM-sub with Iterative Offline Correction.
@@
Prompt 3 Iteration 1 31.89 9.78 84.16 10.58
Iteration 2 36.15 13.22 87.89 9.46
Iteration 3 37.03 13.94 88.29 9.23
Iteration 4 37.11 13.95 88.34 9.23
@@ Appendix C Effect of Perturbation Method
Iterative Prompting Iterative Offline Correction
T-F1↑ G-F1↑ G-BS↑ GED↓ T-F1↑ G-F1↑ G-BS↑ GED↓
Single Verifier Base 15.11 7.72 83.63 12.91 15.11 7.72 83.63 12.91
Iteration 1 19.55 8.78 85.90 11.98 18.97 8.72 86.47 12.10
Iteration 2 21.57 9.33 86.59 11.56 19.65 9.00 86.76 11.96
Iteration 3 22.49 9.89 87.10 11.37 19.69 9.00 86.77 11.95
Unified Verifier Base 15.11 7.72 83.63 12.91 15.11 7.72 83.63 12.91
Iteration 1 21.40 8.78 86.43 11.69 20.46 8.78 87.18 11.90
Iteration 2 24.37 9.50 87.50 11.12 21.54 9.06 87.56 11.71
Iteration 3 26.06 10.22 87.99 10.83 21.57 9.06 87.57 11.71
@@ Appendix D ChatGPT vs GPT-3
Appendix E G2T Results on GenWiki-HIQ
BLEU↑ METEOR↑ TER↓ BERTScore↑
GenWikiFINE-f 35.71 36.67 65.19 93.74
GenWiki-HIQ 48.17 42.03 41.94 95.44

### v3

@@ Abstract
Large language models (LLMs) have shown great abilities of solving various natural language tasks in different domains. Due to the training objective of LLMs and their pre-training data, LLMs are not very well equipped for tasks involving structured data generation. We propose a framework, Prompting with Iterative Verification (PiVe), to improve graph-based generative capability of LLMs. We show how a small language model could be trained to act as a verifier module for the output of an LLM (i.e., ChatGPT, GPT-4), and to iteratively improve its performance via fine-grained corrective instructions. We also show how the verifier module could apply iterative corrections offline for a more cost-effective solution to the text-to-graph generation task. Experiments on three graph-based datasets show consistent improvement gained via PiVe. Additionally, we create GenWiki-HIQ and highlight that the verifier module can be used as a data augmentation tool to help improve the quality of automatically generated parallel text-graph datasets. [^1] Our code and data are available at https://github.com/Jiuzhouh/PiVe.
@@ 1 Introduction
A semantic graph, as a form of graph-structured data, stores information in a machine-accessible way (van Harmelen et al., 2008). Generating a semantic graph from text is known as text-to-graph (T2G) generation and is previously attempted mostly by fine-tuning small language models (Xu et al., 2022; Guo et al., 2020). However, generating graph-structured data remains a challenge for LLMs even in the presence of reasonable number of few-shot examples. In fact, regardless of the number of few-shot examples or prompting style the outputs from LLMs (e.g., GPT-3.5) still contain errors and require correction (§4.4).
@@ 2 Basic Definitions
To prompt the LLM, we use few-shot by showing an example of T2G in the prompt to specify the expected format of the semantic graph (i.e., set of triples). We report experiments under various number of shots (§4.6). The basic form of instruction we use in the prompt is "Transform the text into a semantic graph." followed by a text and a semantic graph pair as a demonstration. We also show results under different prompting strategies (§4.7). Different demonstrations are used for different datasets to adapt to the style of different datasets. We show the used demonstrations in Appendix G.
@@ 3.2 Iterative Prompting
As with the prompt used in the first iteration, we still provide an example in the new prompt for the subsequent iterations. The new prompt is "Transform the text into a semantic graph and also add the given triples to the generated semantic graph." In addition to the text, we also include some triples predicted by the verifier module which LLMs are likely to miss. This explicitly instructs the LLM to generate semantic graph and include the given triples. The given triple set contains the predicted missing triples from each iteration, which prevents the LLM from making the same mistakes as in previous iterations. See Appendix G for the used demonstrations.
@@ 4.2 The LLM and Verifier Modules
ChatGPT (gpt-3.5-turbo) is used as our default LLM to perform the T2G task. [^3] We also compare the graph-based generative capability between ChatGPT and GPT-3 in Appendix D and report the fine-tuned results on small language model in Appendix E. We also experiment with GPT-4 in Subsection 4.7. For verifier module, we use T5-Large (Raffel et al., 2020), and Flan-T5-XXL (Chung et al., 2022) as the backbone models for dataset-specific verifier module, and unified verifier module, respectively. T5 models follow the encoder-decoder architecture and treat all NLP tasks as unified text-to-text transduction tasks. Flan-T5 is instruction-fine-tuned version of T5 which was trained on 1,836 NLP tasks initialized from fine-tuned T5 checkpoint. For T5-large, we fine-tune all parameters for separate verifier modules per each dataset. While for Flan-T5-XXL, we use LoRA (Hu et al., 2022) as a parameter-efficient fine-tuning method, to train a unified verifier module which can follow the instruction. When using the unified verifier, we specify the dataset name in the instructions as datasets have different naming convention for relations.
The verifier is implemented using Pytorch (Paszke et al., 2019) and Transformers (Wolf et al., 2020). For the training, we use Adam optimizer (Kingma and Ba, 2015). Details about hyperparameter setting is provided in Appendix A. For the implementation of parameter efficient training method used in Flan-T5-XXL, we use PEFT (Mangrulkar et al., 2022) and 8-bit quantization technique (Dettmers et al., 2022). All training was done using a single A40 GPU with 48GB of RAM.
@@ 4.3 Evaluation Metrics
Time T-F1↑ G-F1↑ G-BS↑ GED↓ Time T-F1↑ G-F1↑ G-BS↑ GED↓
Single Verifier Base 36.0 13.50 4.89 83.92 13.20 36.0 13.50 4.89 83.92 13.20
Iteration 1 +13.5 17.92 5.78 85.91 12.37 +4.5 17.76 5.83 86.42 12.37
Iteration 2 +4.7 19.46 6.44 86.57 12.08 +1.1 18.51 6.11 86.91 12.19
Iteration 3 +2.1 20.17 6.61 86.83 11.95 +0.2 18.55 6.17 86.94 12.18
Unified Verifier Base 36.0 13.50 4.89 83.92 13.20 36.0 13.50 4.89 83.92 13.20
Iteration 1 +118.9 19.64 6.00 86.39 12.08 +105.0 16.99 5.67 87.08 12.95
Iteration 2 +46.8 22.11 6.44 87.31 11.68 +40.6 17.76 5.67 87.48 12.96
Iteration 3 +21.1 23.11 7.50 87.70 11.35 +10.4 17.85 5.67 87.52 12.96
Table 3: Comparison between Iterative Prompting and Iterative Offline Correction on KELM-sub dataset across all metrics using Single Verifier and Unified Verifier. Time denotes the total inference time in minutes.
@@ 4.4 Results
On WebNLG and GenWiki datasets, single verifier module performs slightly better than unified verifier module. While on KELM-sub dataset, unified module performs far better. We speculate this is due to the size of training data for KELM-sub verifier module being larger than that for WebNLG and GenWiki (as shown in Table 1). Since unified verifier module combines the training data of different datasets, more training data leads to better performance for instruction-tuning. We conducted human evaluation which we include in Appendix F due to page limit.
@@
Instead of iteratively prompting the LLM, another way to utilise the results from verifier module is to append the predicted missing triples to the previously generated graph. The results of the comparison between iterative prompting and iterative offline correction using single verifier module and unified verifier module on KELM dataset is shown in Table 3. Iterative offline correction performs worse than iteratively prompting. This might be because iteratively prompting has the chance of doing self-correction. In each iteration, when we prompt the LLM, the generated graphs can probably correct the mistakes that were made in previous iteration. For example, in Figure 4, in Base the predicted relation regarding birth date is “birth year”, while the reference is “date of birth”. As the PiVe iteration continues, in Iteration 2, the relation “birth year” is regenerated as “date of birth” even though we didn’t mention this in the prompt. Due to the page limit, we report the comparison results on WebNLG and GenWiki datasets in Appendix B. Similarly, iterative prompting can achieve better results than iterative offline correction over all using different verifier modules.
@@ 4.6 Impact of More Shots
While in our main experiments, for cost reason, we used only one-shot demonstrations for the LLM prompting (i.e., GPT-3.5), we show that PiVe is effective in improving the results regardless of the underlying number of shots. Here we report the results of k-shot (k=6, 8, 10) with the iterative offline correction (i.e., only using the LLM once to get the initial graph, while the correction steps are all applied step-by-step and offline). Figure 2 demonstrates the results on KELM-sub using unified verifier with iterative offline correction. The results show, as expected, that PiVe still provides consistent improvement even with the increase in the number of shots. Additionally, as the shots grow the improvement from PiVe also increases.
@@ 4.7 Baselines on LLMs
We also compare PiVe with the recent Self-Refine (Madaan et al., 2023) method, which leverages LLM itself to provide feedbacks for self-refinement. Table 4 shows the self-refine results on KELM-sub dataset. The results show that self-refine could not provide effective feedback, thus leading to the performance drop as the iteration goes. The performance gap is obvious comparing with our PiVe result. Since LLMs are not trained as rigorously on structured data compared to text, expecting them to provide meaningful feedbacks on their outputs is expected to fail.
T-F1↑ G-F1↑ G-BS↑ GED↓
Base 13.50 4.89 83.92 13.20
Iteration 1 14.54 3.00 84.85 13.90
Iteration 2 14.33 2.44 84.57 14.43
Iteration 3 13.79 2.28 84.20 15.18
Table 4: Self-Refine (Madaan et al., 2023) results on KELM-sub using Single Verifier.
Figure 3: Results of using 3 diverse prompts with 6-shot on KELM-sub with Iterative Offline Correction on ChatGPT and GPT-4. The colors represent Base, and corrective iterations 1, 2, 3.
@@ 4.8 Computational Cost and Trade-off
The training and inference of both single verifiers and unified verifier are on a single A40 GPU. Each single verifier takes around 6 hours and the unified verifier takes around 40 hours to train. The computation cost for training of verifiers is a feasible one-off cost. Once the training is finished, the inference of the verification of each instance takes 0.15s for single verifier, and 3.5s for unified verifier. See the Time column in Table 3. Different verifiers present performance-speed trade-offs and are significantly effective in augmenting the LLMs.
@@
4.9 PiVe Examples
In Figure 4, we demonstrate an example from KELM-sub test set using unified verifier. In Base, based on the prediction from the LLM, the verifier module predicts the missing triple [‘‘Francisco Uranga’’, ‘‘occupation’’, ‘‘swimmer’’]. By suggesting this missing triple in the next iteration of prompt, the prediction from LLM includes it. Then, the verifier predicts the missing triple [‘‘Francisco Uranga’’, ‘‘sex or gender’’, ‘‘male’’]. In Iteration 2, both of these two missing triples are included in the prediction from LLM, and at this time, the verifier predicts “Correct”. The prediction from Iteration 2 contains all information correctly in the reference. See another example in Appendix H.
@@ 5 GenWiki-HIQ
Table 5: Results of iterative augmentation on filtered 110K text-graph pairs from GenWikiFINE across four metrics after each iteration. We take the text-graph pairs from Iteration 3 as the created GenWiki-HIQ dataset.
@@
We used the dataset-specific verifier module to do the data augmentation. We conducted four iterations and the evaluation results are shown in Table 5. The results in Base represent the scores over non-parallel graph-text pairs from GenWikiFINE, which have low overlap between graph and text. By using verifier module iteratively, we add more missing triples to the original graph, thus leading the higher quality scores. As the iteration progresses, fewer missing triples are added and we take the augmented graph-text pairs from the last iteration as the final created GenWiki-HIQ dataset. We also conducted G2T experiments in Appendix 5.1 to further demonstrate the quality of GenWiki-HIQ. The G2T model trained on GenWiki-HIQ performs far better than the G2T model trained on GenWikiFINE-f on the human annotated GenWiki test set. This indicates that GenWiki-HIQ contains parallel text-graph pairs with high overlap.
@@
5.1 G2T Results on GenWiki-HIQ
BLEU↑ METEOR↑ TER↓ BERTScore↑
GenWikiFINE-f 35.71 36.67 65.19 93.74
GenWiki-HIQ 48.17 42.03 41.94 95.44
@@ 6.3 Verifiers
Leveraging small models could further improve the performance of LLMs. Cobbe et al. (2021) proposed to solve math word problem by utilising verifier. The verifier is used to judge the correctness of model-generated solutions. During test time, based on multiple candidate solutions generated, verifier calculates the correctness probability and the final answer will be selected by the verifier from the ranked list. Welleck et al. (2023) proposed self-correction, an approach that trains a small model to iteratively apply self-correction. The idea of self-correction looks similar to our PiVe. While Welleck et al. (2023) focuses on the design of a self-correcting language model, PiVe presents a very simple verifier module design and a simple data perturbation strategy to train such model. The ideas presented in our work are developed concurrently and independently.
@@ Appendix B Additional Experiment Result
KELM-sub 58.45 47.26 94.12 8.48
WebNLG 54.77 45.31 93.51 9.11
GenWiki 36.34 29.69 91.14 9.74
Iterative Prompting Iterative Offline Correction
T-F1↑ G-F1↑ G-BS↑ GED↓ T-F1↑ G-F1↑ G-BS↑ GED↓
Single Verifier Base 15.11 7.72 83.63 12.91 15.11 7.72 83.63 12.91
Iteration 1 19.55 8.78 85.90 11.98 18.97 8.72 86.47 12.10
Iteration 2 21.57 9.33 86.59 11.56 19.65 9.00 86.76 11.96
Iteration 3 22.49 9.89 87.10 11.37 19.69 9.00 86.77 11.95
Unified Verifier Base 15.11 7.72 83.63 12.91 15.11 7.72 83.63 12.91
Iteration 1 21.40 8.78 86.43 11.69 20.46 8.78 87.18 11.90
Iteration 2 24.37 9.50 87.50 11.12 21.54 9.06 87.56 11.71
Iteration 3 26.06 10.22 87.99 10.83 21.57 9.06 87.57 11.71
@@
Prompt 3 Base 31.89 9.78 84.16 10.58
Iteration 1 36.15 13.22 87.89 9.46
Iteration 2 37.03 13.94 88.29 9.23
Iteration 3 37.11 13.95 88.34 9.23
@@ Appendix D ChatGPT vs GPT-3
Appendix E Comparison with Fine-tuned Baselines
While our work focuses on the fundamental question of "How can we improve the generative capabilities of black box LLMs in graph generation?", for completeness we also provide results of fine-tuned T5 Raffel et al. (2020) in Table 12. As expected, fine-tuning on large amount of data surpasses few-shot prompting. This underscores the struggle LLMs face in transduction problems, and the need for additional mechanisms (like PiVe) to help LLM improvement.
Appendix F Human Evaluation
After annotation, we took majority voting over the result of each instance, then calculated the number of wins for ChatGPT with or without PiVe. The results are shown in Table 16. From the results, we can see ChatGPT with PiVe wins on 85 out of 105 samples and the total winning rate is over 80%. This indicates the PiVe can effectively improve the graph-based generative capability of LLMs.
For the cases that PiVe failed to improve, we did error analysis and found that there were mainly two types of mistakes that PiVe made: redundancy and inaccuracy. In Figure 6, we demonstrate two examples containing these two types of mistakes shown in red text. In the first example, the triple [‘‘Train song Mermaid’’, ‘‘instrument’’,‘‘Singing’’] predicted by PiVe is redundant. In the second example, the relation ‘‘date Of Retirement’’ in the triple [‘‘Alan Shepard’’,‘‘date Of Retirement",‘‘1963’’] is inaccurate. We speculate these behaviours were caused due to the presence of many similar texts with similar graphs in the training data. During training, PiVe learned the potential connections between these similar graphs, thus leading to redundant and inaccurate triples at prediction.
Appendix G Demonstrations in Prompt
Appendix H PiVe Examples
Figure 9 illustrates another example of PiVe from WebNLG test set using single verification module. In Base, the verification module predict the missing triple [‘‘Agremiação Sportiva Arapiraquense’’, ‘‘ground’’, ‘‘Estádio Municipal Coaracy da Mata Fonseca’’], even though there is a similar triple but containing mistakes in the prediction from the LLM. In Iteration 1, the LLM corrects the mistakes in the previous iteration, and also includes the predicted missing triple. Based on the prediction from the LLM, the verification module predict the missing triple [‘‘Campeonato Brasileiro Série C’’, ‘‘country’’, ‘‘Brazil’’]. In Iteration 2, the verification module predict “Correct” to the final prediction from the LLM. After three iterations using PiVe, the predicted graph contains all information in the reference.